Problem 13-15

The owner of a large machine shop has just finished its financial analysis from the prior fiscal year. Following is an excerpt from the final report:

Net revenue

$

365,000

Cost of goods sold

316,000

Value of production materials on-hand

42,500

Value of work-in-process inventory

41,000

Value of finished goods on-hand

14,500

  1. Compute the inventory turnover ratio (ITR) (Round answer 1 decimal place)

Inventory Turnover Ration ______ per year

  1. Compute the weeks supply (WS) (Don’t round intermediate calculations, round 1 decimal place)

Weeks of supply _____
Answer #1

Answer: Inventory turnover ratio = COGS / Average inventory

Since opening and closing inventory is not given, we'll take total inventory as Average inventory

A) ITR = 316000/ (42500+41000+14500) =3.22

B) Weeks of supply = 52/ITR = 52/3.2244= 16.13 week